OpenStudy (anonymous):

How can I solve x^2-6x-3=4 by completing the square?

OpenStudy (wolf1728):

Move the non-x term to the right x^2 -6x = 7 Take coefficient of x which is -6 divide it by 2 then square it and this becomes 9 Add 9 to both sides of equation. x^2 -6x +9 = 16 Take square root of both sides (x-3) = 4 x = 7